OpenStudy (anonymous):

Select one of the factors of 5x2 + 7x + 2. (5x - 2) (x + 2) (5x + 1) None of the above I don't want the answer but to know how to solve it. Please and thank you :)

OpenStudy (neer2890):

For quadratic equation \[ax ^{2}+bx+c=0\] roots are \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (neer2890):

alternatively \[5x ^{2}+7x+2=0\] \[5x ^{2}+5x+2x+2=0\] 5x(x+1)+2(x+1)=0 (5x+2)(x+1)=0 so (5x+2) is a root of this equation.

OpenStudy (anonymous):

What happened to 7x?

OpenStudy (anonymous):

Ohh, ok. I get it. Thank you

OpenStudy (neer2890):

you're welcome...:)